OpenStudy (anonymous):

Passing through (-2,1) and parallel to the line whose equation is 2x-3y=7. What is the slope form?

OpenStudy (anonymous):

i got for point slope y-1=-2/3(x+2). but i keep gettin the slope form wrong. What is the slope form for this equation.

OpenStudy (anonymous):

the slope is 2/3

OpenStudy (anonymous):

yes, the slope should be 2/3 as ^

OpenStudy (anonymous):

for parallel lines the gradient is the same, so yo have an equation 2x-3y=7, making y the subject of the formula, y=(2/3)x -7, the slope of this line is 2/3 meaning that a line which parallel to that has a same slope

OpenStudy (anonymous):

thanks